OpenStudy (anonymous):

A .50kg (#1) cart is pulled with a 1.0 Newton force for 1.0 second; another cart (#2) with the same mass is pulled with a 2.0 Newton force for .50 seconds. Which cart has the greatest accleration? Explain. a. Which cart has the greatest Impulse? Explain. b. Which cart has the greatest change in momentum? Explain.

OpenStudy (anonymous):

i got for part a.), Cart 2 because it has a larger force and Cart #2 also has the greatest accleration because it has a greater force as well :)

OpenStudy (anonymous):

Are we doing all your homework here? ;) Use F = ma for first part , no times needed. a) impulse = F x time b) impulse IS the change in momentum as: Impulse = ma x time Impulse = m(change in v/t) x t Impulse = m x change in v = change in momentum!

OpenStudy (anonymous):

what would be the acceleration in the first part?

OpenStudy (anonymous):

a=.5m/s?

OpenStudy (anonymous):

It is cart #1 for all 3, yes. To find out the numbers, use the F = ma formula Cart #1: a = F / m = 1 / 0.5 Cart #2: a = F / m = 2 / 0.5

OpenStudy (anonymous):

but cart #1 is 2 and cart #2 is 4 when i calculated so shouldn't it be cart #2 for all of dem?

OpenStudy (anonymous):

sorry yes!! completely. that is excellent, showing you are following it (I am answering ther questions too.. (no excuse)).

OpenStudy (anonymous):

so i use F=ma for all 3 to prove my answer? ^^

OpenStudy (anonymous):

ok so what would be the acceleration

OpenStudy (anonymous):

Just the first part (the acceleration). part a) and b) are linked, they use the impulse formula (look above)

OpenStudy (anonymous):

yeah so wats theacceleration? :)

OpenStudy (anonymous):

look in the post where i said the wrong cart, i wrote it out for you!!

OpenStudy (anonymous):

and you already worked them out..

OpenStudy (anonymous):

wait so whats the acceleration for both carts???

OpenStudy (anonymous):

so i can finish the equation F=ma

OpenStudy (anonymous):

cart 1 is a = 2m/s/s, cart 2 is a = 4 m/s/s

OpenStudy (anonymous):

wait so for part a.) and b.) it is both carts because I= 1x 1= 1 I2= 2 x .5= 1 as well, and since u said Impulse is the same as momentum then both carts have same Impulse as well as same momentums :)

OpenStudy (anonymous):

RIGHT? :D